OpenStudy (anonymous):

Can someone make sure I did this correctly, I'm a little rust on dose calculations. Thanks A robust first grader goes for swimming lessons three times a week, and he enjoys it thoroughly. During the first week of December, he develops an upper respiratory infection, starts coughing continuously, and has a high fever of 39°C. His condition remains the same even after five days of bed rest and conservative measures. Eventually, his mother decides to take him to the doctor who diagnoses it as an acute case of middle ear infection (otitis media), and prescribes amoxicillin for his treatment. He

OpenStudy (anonymous):

weighs 44lb, and the daily dose of amoxicillin for the treatment of the ear infection is 30mg/kg per day in divided doses every 12 hours. • Determine the patient’s body temperature in Fahrenheit. 39 * 1.8 = 70.2 + 32 = 102.2 degrees F • Determine his dosage of amoxicillin in mg. per single dose. To convert from lbs to kg 2.2 lb : 1kg :: 44lb : x kg 2.2x = 44 X = 44 / 2.2 X = 20 kg Weight is 20 kg Recommended 20 kg * 30 mg/kg/24divided q12h = 600mg/24h 600mg/24h 2 doses/24h Single dose is 300mg/dose

OpenStudy (anonymous):

F degree = 9 C/ 5 + 32 = (9 * 39 / 5 ) + 32 = 102.2 F Weigh ( kg) = 44 lb / 2.2 = 20 kg Dose/ 24 hrs = 20 kg * 30 mg * 2 = 600 mg -> Dose/ 12 hrs = 600/2 = 300 mg

OpenStudy (anonymous):

You're absolutely correct :)

OpenStudy (anonymous):

Thank you so much, I thought I was but wanted to make sure :)

OpenStudy (anonymous):

Just make it neater, then it'll be less confused later :)